OpenStudy (vrajrpatel):

Holly deposited $10,000 in an investment account. After 6 years the account is worth approximately $13,400. Which formula represents the value of the account after t years?

OpenStudy (retireed):

Calculate the rate r .... A = P e^rt A/P = e^rt ln(A/P) = rt ln(A/P) / t = r ln(13400/10000) / 6 = r Then plug the variables in to calculate the amount for any t A = 10000 e^rt A = 10000 e^(ln(134/100) / 6 ) t